OpenStudy (firejay5):

Let (x1, y1) & (x2, y2) represent the coordinates of 2 points. Write a formula for finding the coordinates of midpoint M of the segment connecting those two points. Medal will be rewarded for your patronage! :D

OpenStudy (anonymous):

\[\left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right)\] the average in each coordinate